# Statistical Inference with Limited Memory: A Survey

### v1

@@ 3.2 Inference Under Communication Constraints
The topic of inference under communication constraints is not covered in this survey. Nevertheless, as it is closely related to inference under memory constraints (as we will later see), we provide a short and non-exhaustive overview of this problem. The study of statistical inference under communication constrains has originally been considered in the information theoretic literature. Parameter estimation under communication constraints was initially studied by Zhang and Berger [33], who provided a rate-distortion upper bound on the quadratic error in distributed estimation of a scalar parameter using one-way communication under a rate constraint of bits per sample. There has been much follow-up work on this type of problems, especially in the discrete alphabet case, see in particular [34, 35, 36, 37, 38] and references therein. The closely related problem of distributed hypothesis testing against independence under communication constraints was originally treated by Ahlswede and Csisár [39], who provided an exact asymptotic characterization of the optimal tradeoff between the rate of one-way protocols, and the Type I error exponent attained under a vanishing Type II error probability. This result has been extended in many interesting ways, see, e.g., [40, 41, 42, 43, 44, 45, 46]. Specifically, a generalization to the interactive setup with a finite number of rounds was addressed in [47], and the associated one-way communication complexity problem was recently studied via hypercontractivity in [48].
@@
In a different line of work [49, 50], Hadar et al. assumed that the agents have access to an arbitrarily large number of local samples, and are only limited by the number of bits they can exchange. Rather than computing a specific function of finite input sequences with a small error probability under a known distribution (or in the worst case), they assume an unknown distribution and ask to estimate correlations in the minimax sense. They thus define the communication complexity of the inference task to be the minimal number of bits required to guarantee that the minimax risk is below some constant, in the limit of large number of local samples. This definition is different from the definition of communication complexity as presented in the machine learning literature, which is the least possible number of bits needed to be interactively exchanged by Alice and Bob, such that they are able to compute some given function of their local inputs, either exactly or with high probability, see, e.g., [51, 52, 53] .
@@
Finally, there has also been much contemporary interest in distributed inference problems with communication constraints in a different context, where a finite number of i.i.d. samples from an unknown distribution are observed by multiple remotely located parties, who communicate with a data center under a communication budget constraint (one-way or interactively), to obtain an estimate of some property of interest. In these type of problems, the samples observed by the agents are from the same distribution and the regime of interest is typically where the dimension of the problem is high compared to the number of local samples, see, e.g., [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. This problem bares close resemblance to remote source coding under Gaussian noise, known as the CEO problem, which have been extensively studied in the information theory literature, see, for example, [65, 66, 67] and Chapter $33.11$ of Polyanskiy and Wu [68].
4 Algorithm Types
4.1 Time-Varying vs. Time-Invariant
A finite-memory estimation algorithm $\hat{\pi}$ is said to be time-varying, if either the state-transition function $f$, or the state-decision function $g$ (or both) are allowed to depend the time index $n$. Otherwise, namely if both $f$ and $g$ are fixed for all $n$, the algorithm is called time-invariant. We note that any time-varying algorithm can be trivially reduced to a time-invariant one with a $O(\log{n})$ bits overhead, if the total number of available samples $n$ is finite, by simply keeping a clock [^1] We do not account here for the complexity of describing the state transition function $f$. In time-varying algorithms, this description may grow as a function of the number of observed samples.. Thus, in our framework, a time-varying algorithm implicitly means that we assume an external clock is provided “for free”, i.e., without accounting for the memory required to store it. Since what we care about is the memory budget, it seems less preferable to make this kind of an assumption. Moreover, when the number of samples is unbounded, this assumption can lead to absurd results. For example, one can solve the two-armed Bandit problem with only $2$ states [69], and the binary hypothesis testing problem using just $3$ states and with zero error (!) [70, 71]. Finally, as we shall see, some optimal time-invariant algorithms inherently do not keep a clock at all. For all these reasons, in this survey we will restrict our attention to time-invariant algorithms.
4.2 Randomized vs. Deterministic
A randomized algorithm is an algorithm whose state-update and/or state-decision functions depend not only on the samples, but also on some external randomness independent of the samples. A natural question is in what cases randomization even helps? Note that, trivially, there is no loss of generality in assuming that the state-decision function is deterministic; this follows simply since the risk incurred by a random decision rule is some average over risks obtained by deterministic ones, hence once can always pick the risk-minimizing deterministic rule.
5 Methodology
@@ 5.1 Lower Bound Techniques
Reduction to a Communication Complexity Problem
Arguably the earliest known method for deriving lower bounds on the memory complexity of problems dates back to the work of Alon et al. [13] and the problem of finding the memory complexity of approximating the frequency moments in a length $m$ sequence. The authors obtained their lower bounds by reducing the problem to an appropriate communication complexity problem, then they either used some existing result, or proved the required lower bound for the corresponding communication problem. In other words, to find the memory complexity of approximating the $k$th moment $F_{k}$, they lower bound the communication complexity of computing a certain boolean function and reduce it to that of approximating $F_{k}$. The reduction is done by splitting the observed sequence into two parts of equal lengths $x$ and $y$, and allowing two parties, Alice and Bob, to observe the inputs $x$ and $y$, respectively. In this model, Alice and Bob wish to compute the value of a particular boolean function $f(x,y)$ with probability at least $2/3$. The idea is that Alice can run its finite-state algorithm $\mathcal{A}$ on $x$ and transmit the state $s$ of $\mathcal{A}$ to Bob, and then Bob can plug $s$ into his copy of $\mathcal{A}$ and continue the computation on $y$. This computation will return a number $\hat{F}_{k}$ which is an approximation to $F_{k}$ with probability at least $2/3$. If $\hat{F}_{k}$ can be used to determine $f(x,y)$ with error probability less than $1/3$, then the memory used by $\mathcal{A}$ must be at least the one-way communication complexity of computing $f(x,y)$. For example, in one of the lower bounds of [13], the authors reduced the communication complexity of computing equality $EQ(x,y)$ to computing $F_{0}(x\circ y)$. Since the randomized communication complexity of $EQ(x,y)$ is $\asymp\log m$ bits, this established an $\Omega(\log m)$ lower bound for computing $F_{0}$.
@@
The work mentioned above had initiated the research on the memory complexity of approximating statistic of worst case sequences, and many consecutive works have applied the tools of communication complexity to derive lower bounds on different problems, see, e.g., [23, 30, 73, 74, 75]. In the stochastic settings there have also been a few works leveraging communication complexity to attain lower bounds on the memory needed to solve a statistical problem (albeit much less). For example, Woodruff [32] applies this technique for the problem of finding $F_{0}$ when the samples are drawn from a particular distribution, Crouch et al. [76] for the problem of approximating the collision probability and other memory limited problems, Dagan and Shamir [77] for the problem of detecting correlations with little memory and communication, and Dinur [78] for the problem of distinguishing between sampling with or without replacement.
Information Theoretic Lower Bounds
@@
The approach here is to bound the statistical distance between the algorithm output distributions for different values of $\pi$ (the property we want to infer), which usually translates to lower bounding the mutual information between $\pi$ and the memory state. Though there are some examples that leverage this approach directly in the memory-constrained framework (see, e.g., [79, 80, 81]), in general it is quite difficult to quantify the information contraction or find the strong data processing coefficient for many memory-limited problems. Thus, a more simple (yet much weaker) information theoretic approach seeks to derive bounds for a communication constrained problem known as the blackboard model: This is a broadcast model of communication in which each sample is observed by a different player, and each of the players is allowed to communicate at most $b$ bits to all other players. The communication constraints do not allow for arbitrary communication, but rather only “one-pass” protocols in which all information is communicated from left to right. This corresponds to a much less restrictive memory-constrained algorithm whose output at time $t$ is a function of all memory states up to time $t$ (not just the current memory state, as in our model), so the lower bounds for the blackboard problem carry over to our model. In the statistical learning setup, this approach was used to obtain memory-sample trade-offs by Shamir [82], who gave lower bounds on the memory needed for a few learning and estimation problems by reducing them to a canonical ’hide and seek’ problem, and was also used by Steinhardt et al. [83] to obtain lower bounds for memory constrained sparse linear regression.
@@
Any finite states machine that is being fed with an i.i.d. sequence induces a homogeneous Markov chain behaviour. Thus, when dealing with time-invariant algorithms where the time horizon (i.e., number of samples) is sufficiently large with respect to the number of states, we can expect the chain to converge to some limiting distribution, so we can obtain lower bounds by showing that the dependence of that limiting distribution on the property we want to infer is not too strong. This does not hold when the number of states is not much smaller than the number of samples, thus for a time-varying algorithm, in which the number of states grows with time, analyzing the limiting distribution becomes intractable.
Consider, for example, the problem of memory limited binary hypothesis testing, which was solved by Hellman and Cover [84] for randomized algorithms. In their lower bound proof, they define the state likelihood ratio as the ratio between the stationary distributions under both hypotheses. Intuitively, in states with low error probability the state likelihood ratio must be bounded away from $1$. They manged to show that the state likelihood ratios cannot grow too much between adjacent states, and that implied a constraint on the achievable error probability. In a recent work on binary hypothesis testing with memory limited deterministic machines [85], Berg et al. build on probabilistic pigeon-hole arguments that try to capture the fact that different distributions have a non negligible probability of traversing similar Markov trajectories. In particular, they show that for any finite-state machine, there must exist a “bottleneck” state that results in large errors of both types, and is visited with non-negligible probability under both hypotheses.
@@ 5.2 Upper Bound Techniques
Ideally, we would want the Markov chain induced by our FSM driven by the i.i.d. samples to have the following properties:
  • “Good” limiting distribution, e.g., would be highly influenced by the input distribution, so that “small” changes in the distribution property we want to infer would result in “large” changes in the limiting distribution.
  • Fast convergence to the limited/stationary distribution of the chain, that is, good mixing time (corresponding to sample complexity of the algorithm).
Though we are unaware of a general recipe for algorithms that satisfy both properties above, a couple of approaches that were used in the literature may be illuminating; First, when memory is large enough (say, in the neighborhood of $\mathsf{SC}(\delta)\cdot\mathcal{X}$ when taken in bits), an appealing and simple approach to deriving upper bounds is to realize the existing “natural” algorithm from the unconstrained setup in a memory limited scheme. On the other hand, when samples are abundant, the approach of waiting for extreme events was proved useful in formulating order-wise optimal (or close to optimal) algorithms, as we expand on in later sections.
It is worth mentioning that, in some particular problems, essentially only the “natural” or the rare events approach to algorithmic construction needs to be considered. For a concrete example, in the seminal work of Raz [86], the author addressed the problem of parity learning over $\{0,1\}^{d}$ with finite-memory algorithms, and showed that any algorithm for parity learning requires either a quadratic memory size (taken as the number of bits) or an exponential number of samples. Consequently, if one has $O(d^{2})$ memory bits, parity learning can be solved in polynomial time by Gaussian elimination, which is the natural approach. Conversely, if one has an exponential number of samples, parity learning can be solved by waiting for all the singletons, consuming only $d+o(d)$ memory bits, which would be a rare events approach.
@@
In settings where the number of available samples is arbitrarily large, waiting has no cost, thus it makes sense for an agent to focus on capturing suitable rare events in the sample space such that, conditioned on such events, the inference task becomes easier. The memory size is then related to the rarity of the event, since the agent needs to have sufficiently large memory in order to detect it in the first place, but describing a few rare events is usually very light in memory. To exemplify this, consider again the case of binary hypothesis testing. This type of approach has been originally suggested by Cover [70], and studied by others since then. In the construction of Hellman and Cover [84], the finite state machine only changes a state upon observing a maximal or minimal likelihood ratio event. When it reaches an extreme state it also waits for an extreme event, but then exits the state with small probability, due to the use of artificial randomization. In [85], the deterministic binary hypothesis testing scheme is based on observing a long consecutive run of 00’s or a long consecutive run of $1$’s, which have exponentially small probabilities, but conditioned on these events, one hypothesis becomes extremely more likely than its counterpart. Specifically, both waiting for rare events schemes described above attain a risk of $2^{-\Omega(S)}$. This should be juxtaposed with using the natural statistic for this problem, the number of $1$’s in the sample, which leads to poor performance; the quadratic risk cannot decay faster than $2^{-\Omega(\sqrt{S})}$. This follows since in order to count the number of $1$’s in a stream of length $k$ we must keep a clock that counts to $k$, thus overall the number of states used is $S=O(k^{2})$. As it turns out, the deterministic binary hypothesis testing machine that seeks long repetitions can be used as a building block to construct optimal machines in other settings (see [87, 88]), thus it seems that the rare event approach is fundamental for finite memory problems where the number of samples is very large.
@@ 6 Simple Hypothesis Testing
One of the earliest appearances of a memory constraint problem in the statistical inference literature is in the hypothesis testing context. Recall that in this problem, i.i.d. samples are drawn an unknown distribution $P\in\mathcal{F}=\{P_{0},\ldots,P_{d-1}\}$, where under hypothesis $\mathcal{H}_{i}$ we have $P=P_{i}$ with $\pi(P_{i})=i$, and we want to guess the correct underlying hypothesis, under the $0/1$ loss, $\ell(\pi,\hat{\pi})=\operatorname{\mathds{1}}(\pi\neq\hat{\pi})$. For the majority of this section we will be interested in the binary hypothesis testing problem, in which =$\mathcal{F}=\{P_{0},P_{1}\}$. We show that even this simple version of maybe the most basic inference task is still difficult under memory limitations. In most of the section, we will be interested in the limit of the problem when $n\rightarrow\infty$, thus following Eq. (12), we define the asymptotic probability of error of an $S$-state algorithm, when the true hypothesis is $\mathcal{H}_{i}$, to be
@@
Interest in the limited memory hypothesis testing problem seems to have been sparked by the work of Robbins [1] on the Two-Armed Bandit problem. In this setup, a player is given two coins, with Bernoulli parameters $p_{1}$ and $p_{2}$ respectively, whose values are unknown. The player is required to maximize the long-run proportions of “heads” obtained, by successively choosing which coin to flip at any moment. Robbins proposed an algorithm that works with a limited memory of $S$ states and attains the optimal performance as $S$ approaches infinity. Robbin’s algorithm was later improved by Isbell [93], Smith and Pike [94], and Samuels [95], where the latter improvement arises from a randomized algorithm.
Departing from the restriction to time-invariant algorithms imposed in the works mentioned above, Cover [69] discovered a time-varying finite memory algorithm that achieves the optimal proportion $\max\{p_{1},p_{2}\}$, asymptotically in the number of tosses, with $S=2$. Essentially, Cover suggested to modify the problem by allowing to keep a clock, such that at each stage we know how many tosses we have made so far. Cover’s algorithm interleaves test blocks and trial blocks, where a test block is a sequence of tosses that test the hypothesis $p_{1}>p_{2}$ vs. $p_{2}>p_{1}$, and a trial block is a sequence of exclusive tosses of the “favorite” coin resulting from the preceding test block. The choice of the sequence of test-trial block sizes then assures that the maximum is attained. The intuitions and tools developed for the Two-Armed Bandit problem were used by Cover in a subsequent paper addressing the binary hypothesis problem [70], in which he described a time-varying finite memory machine that solves the binary hypothesis testing problem with $S=4$ states (this was later improved to $S=3$ by Koplowitz [71], who more generally showed that for $M$-ary hypothesis testing with a time-varying finite memory, $M+1$ states are necessary and sufficient to resolve the correct hypothesis with an error probability that approaches zero asymptotically). Specifically, Cover describes an algorithm with four memory states that, given samples from a $\mathsf{Bern}(p)$ distribution, can decide between the hypothesis $\mathcal{H}_{0}:p=p_{0}$ and $\mathcal{H}_{1}:p=p_{1}>p_{0}$ with error probability that converges to zero almost surely. [^2] We note that the result in the paper is stronger, solving a composite hypothesis testing problem of the form $\mathcal{H}_{0}:p>p_{0}$ vs. $\mathcal{H}_{1}:p<p_{0}$.

### v2

@@ 4.2 Inference Under Communication Constraints
The topic of inference under communication constraints is not covered in this survey. Nevertheless, as it is closely related to inference under memory constraints (as we will later see), we provide a short and non-exhaustive overview of this problem. The study of statistical inference under communication constrains has originally been considered in the information theoretic literature. Parameter estimation under communication constraints was initially studied by Zhang and Berger [41], who provided a rate-distortion upper bound on the quadratic error in distributed estimation of a scalar parameter using one-way communication under a rate constraint of bits per sample. There has been much follow-up work on this type of problems, especially in the discrete alphabet case, see, in particular [42, 43, 44, 45, 46] and references therein. The closely related problem of distributed hypothesis testing against independence under communication constraints was originally treated by Ahlswede and Csisźar [47], who provided an exact asymptotic characterization of the optimal tradeoff between the rate of one-way protocols, and the Type I error exponent attained under a vanishing Type II error probability. This result has been extended in many interesting ways, see, e.g., [48, 49, 50, 51, 52, 53, 54]. Specifically, a generalization to the interactive setup with a finite number of rounds was addressed in [55], and the associated one-way communication complexity problem was recently studied via hypercontractivity in [56].
@@
In a different line of work [57, 58], Hadar et al. assumed that the agents have access to an arbitrarily large number of local samples, and are only limited by the number of bits they can exchange. In this setup, it is assumed that the underlying distribution is unknown, and the goal is to estimate correlations in the minimax sense. They thus define the communication complexity of the inference task to be the minimal number of bits required to guarantee that the minimax risk is below some constant, in the limit of large number of local samples. This definition is different from the definition of communication complexity as presented in the machine learning literature, which is the least possible number of bits needed to be interactively exchanged by Alice and Bob, such that they are able to compute some given function of their local inputs, either exactly or with high probability, see, e.g., [59, 60, 61] .
@@
Finally, there has also been much contemporary interest in distributed inference problems with communication constraints in a different context, where a finite number of i.i.d. samples from an unknown distribution are observed by multiple remotely located parties, who communicate with a data center under a communication budget constraint (one-way or interactively), to obtain an estimate of some property of interest. In these type of problems, the samples observed by the agents are from the same distribution and the regime of interest is typically where the dimension of the problem is high compared to the number of local samples, see, e.g., [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 3, 72]. This problem bares close resemblance to remote source coding under Gaussian noise, known as the CEO problem, which have been extensively studied in the information theory literature, see, for example, [73, 74, 75] and Chapter $33.11$ of Polyanskiy and Wu [76].
5 Methodologies
In this section, we provide an overview of common methodologies and ideas used in deriving performance bounds in finite-memory inference. For lower bounds, we describe the method of reduction to a communication complexity problem (which originated from works on the streaming model), the information theoretic approach, and the approach of Markov chain analysis. We then delineate some useful guidelines for upper bound derivations.
@@ 5.1 Lower Bound Techniques
Reduction to Communication Complexity
It is often possible to reduce a finite memory inference problem to a communication complexity problem, essentially replacing the bounded memory constraint by a bounded communication constraint. The high level idea is as follows: Suppose we have an $S$-state algorithm that can solve the inference problem given $n$ samples, within the desired level of risk. Now, we can partition the samples into two parts $a$ and $b$, giving $a$ to Alice and $b$ to Bob. Alice can run the algorithm on $a$, and send the final state to Bob using $\log{S}$ bits. Bob can then start his machine at Alice’s last state, and continue running it on $b$, thereby emulating the entire inference procedure. One can now try to find some function $f(a,b)$ such that solving the inference problem also allows Bob to compute $f(a,b)$ with good probability. The communication complexity of reliably computing $f(a,b)$, i.e., the minimal number of bits Alice and Bob need to exchange to that end, is therefore a lower bound on $\log{S}$, and can sometimes be evaluated directly. We note that this useful reduction originates from the streaming literature, particularly from the seminal work of Alon et al. [21] on the memory complexity of approximating the frequency moments of a finite sequence, a paper that initiated the research on the memory complexity of data streams statistics approximation.
@@
Following the footsteps of [21], many subsequent works applied the tools of communication complexity to derive lower bounds on different problems in the streaming model, see, e.g., [31, 38, 77, 78, 79]. Applications of this reduction approach in the memory-limited statistical inference literature are few and far between. Among those are the works of Woodruff [40] on counting distinct elements when the samples are drawn from a particular family of distributions; Crouch et al. [80] on approximating the collision probability and other memory limited problems; Dagan and Shamir [81] on detecting correlations with little memory and communication; and Dinur [82] on distinguishing between sampling with or without replacement.
The Information Theoretic Approach
@@
Information-theoretic based lower bounds for the memory constrained inference problem typically rely on lower bounding the mutual information between $\pi$, the property we want to infer, and the memory state (or, similarly, bounding the statistical distance between the algorithm output distributions for different values of $\pi$). Capturing the tradeoff between how large this mutual information can be, and the number of available memory states, can be thought of as a variant of the information bottleneck problem [83, 84]. Though there are some examples that leverage this approach directly (see, e.g., [85, 86, 87]), in general it is quite difficult to quantify the information contraction or find the strong data processing coefficient for many memory-limited problems. Thus, a more simple (yet significantly looser) information theoretic approach seeks to derive bounds for a communication constrained problem known as the blackboard model: This is a broadcast model of communication in which each sample is observed by a different player, and each of the players is allowed to communicate at most some constant number of bits to all other players. The communication constraints do not allow for arbitrary communication, but rather only one-pass protocols in which all information is communicated from left to right. This corresponds to a much less restrictive memory-constrained algorithm whose output at time $t$ is a function of all memory states up to time $t$ (not just the current memory state, as in our model), so the lower bounds for the blackboard problem carry over to our model. In the statistical learning setup, this approach was used to obtain memory-sample trade-offs by Shamir [88], who gave lower bounds on the memory needed for a few learning and estimation problems by reducing them to a canonical ’hide and seek’ problem, and was also used by Steinhardt et al. [89] to obtain lower bounds for memory constrained sparse linear regression.
@@
Another approach is to leverage the fact that any finite state machine driven by an i.i.d. sequence induces a homogeneous Markov chain behaviour. Thus, for time-invariant algorithms where the number of processed samples is sufficiently large w.r.t the number of states, one can expect the chain to converge to some limiting distribution, which allows us to obtain lower bounds by showing that the dependence of that limiting distribution on the property we want to infer is not too strong. We note that this approach breaks down for a time-varying algorithm, since the number of states grows with time and hence analyzing the limiting distribution becomes intractable.
An early example of this method appears in the seminal paper of Hellman and Cover [5] on memory-limited binary hypothesis testing. There, the authors showed that the state likelihood ratio between the stationary distributions under the two hypotheses cannot change too quickly between states, and leveraged this to obtained a lower bound on the achievable error probability. A more recent example is the work of Berg et al. [90] on binary hypothesis testing with memory limited deterministic machines. There, the authors build on probabilistic pigeon-hole arguments that capture the fact that different distributions have a non-negligible probability of traversing similar Markov trajectories. In particular, they showed that for any finite-state machine, there must exist a “bottleneck” state that results in large errors of both types, and is visited with non-negligible probability under both hypotheses.
@@ 5.2 Upper Bound Techniques
Loosely speaking, to construct good upper bound, we want the Markov chain induced by our finite state algorithm to have the following high-level properties:
  • “Good” limiting distribution, e.g., strongly influenced by the input distribution, so that small changes in the property we want to infer would result in large changes in the limiting distribution.
  • Fast convergence to the limiting / stationary distribution, that is, short mixing time. Such behavior will guarantee that a small number of samples suffice for approaching the limiting distribution, which, in turn, results in a small sample complexity.
Though we are unaware of a general recipe for algorithms that satisfies both properties above, a couple of approaches that were used in the literature may be illuminating: First, when memory is large enough (say $\gtrapprox\mathsf{SC}(\delta)\cdot\log|\mathcal{X}|$ in bits), an appealing and simple approach to deriving upper bounds is to realize the existing “natural” algorithm from the unconstrained setup in a memory limited scheme. On the other hand, when samples are abundant, the approach of waiting for extreme events has been useful in formulating order-wise optimal (or close to optimal) algorithms, as we expand on in later sections.
It is worth mentioning that, in some particular problems, essentially only the “natural” or the rare events approach to algorithmic construction needs to be considered. For a concrete example, in the seminal work of Raz [91], the author addressed the problem of parity learning over $\{0,1\}^{d}$ with finite-memory algorithms, and showed that any algorithm for parity learning requires either a quadratic memory size in bits, or an exponential number of samples. Consequently, if one has $O(d^{2})$ memory bits, parity learning can be solved in polynomial time by Gaussian elimination, which is the natural approach. Conversely, if one has an exponential number of samples, parity learning can be solved by waiting for all the singletons, consuming only $d+o(d)$ memory bits, which would be a rare events approach.
@@
In settings where the number of available samples is arbitrarily large, waiting has no cost, thus it makes sense to focus on capturing suitable rare events in the sample space such that, conditioned on such events, the inference task becomes easier. The memory size is then related to the rarity of the event, since we need to have sufficiently large memory in order to detect it in the first place, but describing a few rare events usually does not take much memory. This type of approach has been originally suggested by Cover [4] for the binary hypothesis testing problem, and has been studied by others since then. In the construction of Hellman and Cover [5], the finite state machine only changes a state upon observing a maximal or minimal likelihood ratio event. When it reaches an extreme state it also waits for an extreme event, but then exits the state with small probability, due to the use of artificial randomization. The deterministic binary hypothesis testing scheme of Berg er al. [90] is based on observing a long consecutive run of 00’s or a long consecutive run of $1$’s, which have exponentially small probabilities, but conditioned on these events, one hypothesis becomes extremely more likely than its counterpart. Specifically, both rare events schemes described above attain a risk of $2^{-\Omega(S)}$. This should be juxtaposed with using the natural statistic for this problem, the number of $1$’s in the sample, which leads to poor performance; the quadratic risk cannot decay faster than $2^{-\Omega(\sqrt{S})}$. This follows since in order to count the number of $1$’s in a stream of length $k$ we must keep a clock that counts to $k$, thus overall the number of states used is $S=O(k^{2})$. As it turns out, the deterministic binary hypothesis testing machine that seeks long repetitions can be used as a building block to construct optimal machines in other settings (see [92, 93]), thus it seems that the rare event approach is fundamental for finite memory problems where the number of samples is very large.
@@ 6 Simple Hypothesis Testing
One of the earliest appearances of a memory constraint problem in the statistical inference literature is in the hypothesis testing context. Recall that in this problem, i.i.d. samples are drawn from an unknown distribution $P\in\mathcal{F}=\{P_{0},\ldots,P_{d-1}\}$, where under hypothesis $\mathcal{H}_{i}$ we have $P=P_{i}$, and the property we want to guess is the true underlying hypothesis, $\pi(P_{i})=i$, under the $0/1$ loss, $\ell(\pi,\hat{\pi})=\operatorname{\mathds{1}}(\pi\neq\hat{\pi})$. For the majority of this section we will be interested in the binary hypothesis testing problem, in which =$\mathcal{F}=\{P_{0},P_{1}\}$. We show that even this simple version of maybe the most basic inference task is still difficult under memory limitations. In most of the section, we will be interested in the limit of the problem when $n\rightarrow\infty$, thus following Eq. (12), we define the asymptotic probability of error of an $S$-state algorithm, when the true hypothesis is $\mathcal{H}_{i}$, to be
@@
Interest in the limited memory hypothesis testing problem seems to have been sparked by the work of Robbins [6] on the Two-Armed Bandit problem. In this setup, a player is given two coins, with Bernoulli parameters $p_{1}$ and $p_{2}$ respectively, whose values are unknown. The player is required to maximize the long-run proportions of “heads” obtained, by successively choosing which coin to flip at any moment. Robbins proposed an algorithm that works with a limited memory of $S$ states and attains the optimal performance only as $S$ approaches infinity. Robbin’s algorithm was later improved by Isbell [98], Smith and Pike [99], and Samuels [100], where the latter improvement arises from a randomized algorithm.
Departing from the restriction to time-invariant algorithms imposed in the works mentioned above, Cover [18] discovered a time-varying finite memory algorithm that achieves the optimal proportion $\max\{p_{1},p_{2}\}$, asymptotically in the number of tosses, with $S=2$. Essentially, Cover suggested to modify the problem by allowing to keep a clock, such that at each stage we know how many tosses we have made so far. Cover’s algorithm interleaves test blocks and trial blocks, where a test block is a sequence of tosses that test the hypothesis $p_{1}>p_{2}$ vs. $p_{2}>p_{1}$, and a trial block is a sequence of exclusive tosses of the “favorite” coin resulting from the preceding test block. The choice of the sequence of test-trial block sizes then assures that the maximum is attained. The intuitions and tools developed for the Two-Armed Bandit problem were used by Cover in a subsequent paper addressing the binary hypothesis problem [4], in which he described a time-varying finite memory machine that solves the binary hypothesis testing problem with $S=4$ states (this was later improved to $S=3$ by Koplowitz [19], who more generally showed that for $M$-ary hypothesis testing with time-varying finite memory, $M+1$ states are necessary and sufficient to resolve the correct hypothesis with an error probability that approaches zero asymptotically). Specifically, Cover describes an algorithm with four memory states that, given samples from a $\mathsf{Bern}(p)$ distribution, can decide between the hypothesis $\mathcal{H}_{0}:p=p_{0}$ and $\mathcal{H}_{1}:p=p_{1}>p_{0}$ with error probability that converges to zero almost surely. [^3] We note that the result in the paper is stronger, solving a composite hypothesis testing problem of the form $\mathcal{H}_{0}:p>p_{0}$ vs. $\mathcal{H}_{1}:p<p_{0}$.